# Multimodal ”Puppeteer”: An Exploration of Robot Teleoperation Via Virtual Counterpart with LLM-Driven Voice and Gesture Interaction in Augmented Reality

### v1

@@ V-C3 Summary of Qualitative Feedback
Non-roboticists generally responded positively to the GO interaction, finding it more natural and intuitive, even when encountering technical limitations. One participant remarked, “I prefer the gesture one since the implementation is simple and effective,” while another described the interaction as “great”. However, a consistent theme among this group was difficulty with voice recognition—several users noted that the system had trouble distinguishing commands from environmental sounds, or that the voice interface felt unpredictable. For instance, one participant stated, “For the voice command, it may cause some errors because it might recognizes surrounding sounds,” and another recommended, “If you can use the microphone of the headset, it may reduce misidentification.” These comments suggest that trust in system responsiveness—especially for voice commands—is a central concern for novice users. Their improvement suggestions focused on user comfort and experience, including adding emotional expressiveness through facial interaction or simplifying the voice interface. Notably, their focus leaned toward making the system more playful, accessible, and error-tolerant rather than technically refined.
@@
In summary, the data illustrate a cognitive and experiential divide between the two groups. Non-roboticists favor simplicity, clarity, and responsiveness, often forgiving of technical imperfections if the interface remains accessible. Roboticists, conversely, demand technical precision, stability, and ergonomic viability, expressing frustration when those expectations are unmet. These divergent user needs suggest that future versions of the AR puppeteer system should adopt adaptive interface strategies, offering more guided, simplified interactions for novices, and more fine-tuned, responsive controls for expert users. Additionally, both groups’ repeated emphasis on improving the voice modality signals its current status as a bottleneck for interaction fluidity and trustworthiness.
VI Discussion
@@ VI-A Reflections and Insights
The qualitative data provided important context for interpreting these patterns. Participants frequently described the GO condition as more natural, responsive, and intuitive, reinforcing the strong usability ratings in quantitative metrics. In contrast, VG was perceived as less reliable due to issues with voice recognition, accent sensitivity, and a lack of fluid integration between modalities. While some users acknowledged voice interaction as beneficial in reducing physical effort—especially during repetitive tasks—most reported that the interruptions and mental switching detracted from overall efficiency and satisfaction.
A central concern raised on both conditions was system responsiveness. Participants emphasized the importance of clear, timely feedback and reliable control mapping. This was particularly evident in the VG condition, where inconsistencies in system response undermined user trust and disrupted task continuity. These findings suggest that system predictability and feedback fidelity are critical for maintaining user confidence, especially in real-time HRI scenarios. As such, future development should prioritize seamless integration of modalities, robust input recognition, and consistent feedback loops to support more fluid and trustworthy user experiences.
VI-B Relections from (Non-)Roboticists
@@ VI-C Design Implications
Building on the observed differences between user groups, this study reveals critical considerations for designing multimodal interactions in HRI systems within AR. While multimodality is often associated with increased flexibility and inclusiveness, our findings imply that such benefits are not automatically achieved through input diversity alone. The VG condition, while offered more enjoyable and intuitive experience for non-experts, introduced notable usability challenges, particularly for experienced users, due to synchronization issues between modalities, intermittent system responsiveness, and the cognitive cost of switching input modes. These friction points disrupted interaction flow, undermining the promise of multimodal synergy. As mentioned previously, the effectiveness of multimodal input in AR-HRI loop is highly contingent on user expertise, and without proper contextual integration, voice and gesture inputs can compete rather than complement each other.
@@
Designing multimodal interaction in AR for robotics therefore requires more than just offering multiple input options; it demands a cohesive interaction model that adapts to users’ intent, experience level, and task complexity. Future systems should implement intelligent modality adaptation where the system can hierarchically prioritize or blend modalities based on the contexts and user behaviors. For example, the GO condition could be foregrounded for spatial movements, while voice commands could be contextually triggered for high-level actions or when gestures interactions are not feasible. Furthermore, multimodal systems should feature transparent feedback mechanisms that inform users of current input states and system expectations, reducing ambiguity and improving perceived system resilience. Ultimately, a user-centric, context-aware approach to multimodal design is essential for achieving both performance efficiency and accessibility across varying levels of expertise in designing effective HRI systems within AR.
VI-D Generalizability and Real-World Use
@@
Although conducted in a controlled laboratory environment, the study’s findings have broader implications for deploying AR-mediated robotic puppeteer systems in real-world contexts. The task design, which was centered on spatial manipulation, target matching, and time efficiency, reflects fundamental HRI tasks in domains such as remote inspection, field robotics, warehouse automation or even hazardous environments due to the advantage of interacting with the vritual robot. The clear performance advantage of GO interaction suggests that in high-stakes or time-sensitive scenarios, streamlined unimodal systems may currently offer more reliable and efficient user control. This is particularly relevant for environments where consistency, low latency, and minimal input error are critical.
@@
However, the potential of multimodal systems remains highly relevant for real-world application, especially in situations that demand flexible user input due to physical or cognitive constraints. For instance, in hands-busy contexts such as fieldwork involving protective gear or robotic surgery, voice commands could reduce cognitive and physical strain if implemented reliably. Similarly, in visually cluttered settings, gesture recognition may offer a more dependable alternative. Generalizability, however, depends heavily on system robustness and adaptability. Factors such as fluctuating lighting conditions, limited user training, and unexpected physical interruptions must be considered in future deployment. This underscores the importance of building multimodal systems that are resilient to environmental variability and personalized to user needs. Our findings suggest that while gesture-only systems may serve as a strong foundation, the long-term success of multimodal interaction in real-world HRI will reply on resolving current integration challenges and fostering adaptive, user-aware design paradigms that scale effectively beyond the lab.
VI-E Limitations and Future Work
@@
While the study yielded valuable insights into the comparative effectiveness of GO and VG interactions, several limitations should be acknowledged. One key limitation concerns the technical maturity of the system, particularly the voice command implementation. The VG condition relied on an LLM-driven speech processing module, which, while offering semantic richness, still beard latency, inconsistency in recognition, and a lack of robustness to accents. Participants frequently experienced breakdowns in interaction flow due to delayed or unrecognized voice inputs. These issues likely impacted both subjective workload and task performance. In its current form, the voice interaction pipeline lacks the speed and reliability required for precise instant robotic control tasks, especially those demanding immediate feedback. Additionally, while the gesture interface was generally perceived as intuitive, it suffered from sensitivity issues and occasional misinterpretations, particularly during prolonged or rapid interactions. The gesture recognition system did not always respond fluidly to user input, and its fragility (particularly during fast hand movements) resulted in system interruptions, which disrupted task execution and undermined user trust. Furthermore, the current task design is monotonous, that the voice interaction was only used in the first task stage, which would potentially impact the task performance and user perceptions.
@@
From a design perspective, future work should address the limitations of both input modalities. The voice command techniques can benefit from a more robust and adaptive recognition engine, potentially fine-tuned to specific domain vocabularies or using real-time feedback to clarify ambiguous quaries. For gesture input, improvements in gesture-to-action mapping, more resilient tracking algorithms, and clearer visual or haptic feedback can mitigate usability issues and prevent interruptions during rapid movement. Furthermore, the current usage of modalities in task design lacks variety, which can be improved by designing more combinations of using different modalities in different task stages. Finally, expanding the system’s testing beyond a controlled lab setting and diversifying task scenarios will be essential for evaluating the generalizability to real-world HRI use cases.
VII Conclusion
@@
This study explored the effectiveness of gesture-only versus multimodal (VG) interaction in an AR-based robot puppeteering system. Results showed that GO consistently outperformed VG in task performance, usability, and user satisfaction. While voice input offered some benefits for novice users, it introduced cognitive overhead and system inefficiencies that hindered overall performance. These findings highlight the importance of optimizing modality integration and suggest that hand gesture-based interaction remains a robust solution for robot puppeteer metaphoric system in AR. Future work should explore adaptive, user-aware multimodal systems that balance flexibility with reliability across diverse user contexts.
Acknowledgments
This should be a simple paragraph before the References to thank those individuals and institutions who have supported your work on this article.

### v2

@@ 5.3.3 Summary of Qualitative Feedback
Non-roboticists generally responded positively to the GO interaction, finding it more natural and intuitive, with affirmation on VG interaction. One participant remarked, “I like the gesture one since the implementation is simple and effective,” while another described the interaction as “great”. Even there was some difficulties with voice recognition—several users noted that the system had trouble distinguishing commands from environmental sounds, or that the voice interface felt unpredictable, they praised this modality due to its convenience. For instance, one participant stated, “For the voice command, it may cause some errors because it might recognize surrounding sounds, but it’s so convenient to instruct a robot with a simple voice input besides using gestures” and another recommended, “It didn’t recognize my pronunciation at first try, but it eventually worked and executed the command correctly, which helped processing the task.” These comments suggest that novice user appreciated the multimodality due to it offers another option during interaction.
@@
In summary, the data illustrates a cognitive and experiential divide between the two groups. Non-roboticists favor simplicity, clarity, and responsiveness, often forgiving of technical imperfections if the interface remains accessible. Roboticists, conversely, demand technical precision, stability, and ergonomic viability, expressing frustration when those expectations are unmet. These divergent user needs suggest that future versions of the AR puppeteer system should adopt adaptive interface strategies, offering more guided, simplified interactions for novices, and more fine-tuned, responsive controls for expert users. Additionally, both groups’ repeated emphasis on improving the voice modality signals its current status as a bottleneck for interaction fluidity and trustworthiness.
6 Discussion
@@ 6.1 Reflections and Insights
The qualitative data provided important context for interpreting these patterns. Participants frequently described the GO condition as more natural, responsive, and intuitive, reinforcing the strong usability ratings in quantitative metrics. In contrast, VG was perceived as less reliable due to occasional issues appeared on voice recognition, accent sensitivity, and a lack of fluid integration between modalities. A number of users (more on non-roboticsists’ side) acknowledged voice interaction as beneficial in reducing physical effort – especially during repetitive tasks, however, they also reported that the interruptions and mental switching detracted from overall efficiency and satisfaction.
A central concern raised on both conditions was system responsiveness. Participants emphasized the importance of clear, timely feedback and reliable control mapping. This was particularly evident in the VG condition, where inconsistencies in system response potentially undermined user trust and disrupted task continuity. These findings suggest that system predictability and feedback fidelity are critical for maintaining user confidence, especially in real-time HRI scenarios. As such, future development should prioritize seamless integration of modalities, robust input recognition, and consistent feedback loops to support more fluid and trustworthy user experiences.
6.2 Relections from (Non-)Roboticists
@@ 6.3 Design Implications and Guidelines
Building on the observed differences between user groups, this study reveals critical considerations for designing multimodal interactions in HRI systems within AR. While multimodality is often associated with increased flexibility and inclusiveness, our findings imply that such benefits are not automatically achieved through input diversity alone. The VG condition, while offered more enjoyable and intuitive experience for non-experts, introduced notable usability challenges, particularly for experienced users, due to synchronization issues between modalities, intermittent system responsiveness, and the cognitive cost of switching input modes. As mentioned previously, the effectiveness of multimodal input in AR-HRI loop is highly contingent on user expertise, and without proper contextual integration, voice and gesture inputs can compete rather than complement each other.
@@
Designing multimodal interaction in AR for robotics therefore requires more than just offering multiple input options; it demands a cohesive interaction model that adapts to users’ intent, experience level, and task complexity. Future systems should implement intelligent modality adaptation where the system can hierarchically prioritize or blend modalities based on the contexts and user behaviors. For example, the GO condition could be foregrounded for spatial movements, while voice commands could be contextually triggered for high-level actions or when gestures interactions are not feasible. Furthermore, multimodal systems should feature transparent feedback mechanisms that inform users of current input states and system expectations, reducing ambiguity and improving perceived system resilience. Ultimately, a user-centric, context-aware approach to multimodal design is essential for achieving both performance efficiency and accessibility across varying levels of expertise in designing effective HRI systems within AR. Based on the observed use patterns, we distill five actionable guidelines with some expertise-aware examples for designing multimodal (voice+gesture) AR robot-puppeteering systems as displayed in Table 2.
[…]
G1 Identify domain expertise and set defaults Detect user expertise; start experts in Gesture-Only (GO), novices in Voice+Gesture (VG). First run: experts start in GO with gesture gain and a switch to VG; non-experts start in VG with command overlay (e.g., ”move to XX”, ”pinch and hold”) and a switch to GO.
G2 Allocate roles to modalities Use gesture for continuous, spatial control; voice for discrete, high-level, or hands-busy tasks.
G3 Minimize modalities switching cost Enable seamless switch such as hold-to-talk; keep the context across modes; avoid re-targeting the objects.
G4 Provide transparent feedback mechanism Always show system state; provide clear error or progress cues.
G5 Onboard progressively and adapt interaction over time Begin with more guidance first; gradually reduce prompts and adjust sensitivity as users gain fluency. Experts start with GO; enable adjusting gesture sensitivity and showing/hiding voice settings. Non-experts start with VG with cues (e.g., visual tips); as accuracy improves, cues fade and gestures become primary.
Table 2: Design guidelines for AR-based robot puppeteering with voice+gesture interaction.
6.4 Generalizability and Real-World Use
@@
Although conducted in a controlled laboratory environment, the study’s findings have broader implications for deploying AR-mediated robotic puppeteer systems in real-world contexts. The task design, which was centered on spatial manipulation, target matching, and time efficiency, reflects fundamental HRI tasks in domains such as remote inspection, field robotics, warehouse automation, or operation in hazardous environments where interacting with a virtual robot can reduce risk. The clear performance advantage of GO interaction suggests that in high-stakes or time-sensitive scenarios, streamlined unimodal systems may currently offer more reliable and efficient user control. This is particularly relevant for environments where consistency, low latency, and minimal input error are critical.
From a multimodal interaction perspective, our results highlight a tension between the theoretical benefits of multimodality and the practical costs of integrating multiple channels. Multimodal theory suggests that combining modalities can provide complementary strengths (e.g., gestures for continuous spatial control and speech for discrete, semantic commands), redundancy for error tolerance, and mutual disambiguation when signals are ambiguous. In our setup, however, the additional integration overhead, recognition errors, and temporal misalignment in VG meant that these theoretical benefits did not fully materialize and, in some cases, hindered performance. This underlines that “more modalities” are not inherently better; rather, their value depends on how well they are aligned with task demands, temporal constraints, and users’ cognitive resources.
@@
Nevertheless, the potential of multimodal systems remains highly relevant for real-world applications, especially in situations that demand flexible user input due to physical or cognitive constraints. For instance, in hands-busy contexts such as fieldwork involving protective gear or robotic surgery, voice commands could reduce cognitive and physical strain if implemented reliably. Similarly, in visually cluttered settings, gesture recognition may offer a more dependable alternative. Generalizability, however, depends heavily on system robustness and adaptability. Factors such as fluctuating lighting conditions, limited user training, and unexpected physical interruptions must be considered in future deployment. This underscores the importance of building multimodal systems that are resilient to environmental variability and personalized to user needs. Our findings suggest that while gesture-only systems may serve as a strong foundation, the long-term success of multimodal interaction in real-world HRI will rely on resolving current integration challenges and fostering adaptive, user-aware design paradigms that operationalize multimodal principles in a task- and context-sensitive manner.
6.5 Limitations and Future Work
@@
While the study yielded valuable insights into the comparative effectiveness of GO and VG interactions, several limitations should be acknowledged. One key limitation concerns the technical maturity of the system, particularly the voice command implementation. The VG condition relied on an LLM-driven speech processing module which, although offering semantic richness, still exhibited latency, inconsistency in recognition, and limited robustness to accents. Participants occasionally experienced breakdowns in interaction flow due to delayed or unrecognized voice inputs. These issues likely affected both subjective workload and task performance. In its current early-stage form, the voice interaction pipeline lacks the speed and reliability required for precise, time-critical robotic control tasks, especially those demanding immediate feedback.
Although the gesture interface was generally perceived as intuitive, it also suffered from sensitivity issues and occasional misinterpretations, particularly during prolonged or rapid interactions. The gesture recognition system did not always respond fluidly to user input, and the robot’s inherent fragility (e.g., when users moved their hands quickly) led to safety interruptions that disrupted task execution and undermined user trust. Furthermore, the current task design is relatively monotonous, and voice interaction was only used in the first task stage, which may have influenced both task performance and user perceptions of the VG condition.
A further limitation is the absence of baseline conditions using established teleoperation interfaces such as joysticks, touchscreens, or 6-DoF controllers. As a result, our findings should be interpreted strictly as relative differences between GO and VG within our AR puppeteer context, rather than as evidence that either modality outperforms conventional teleoperation tools. Future work should embed our AR puppeteer interface into a broader comparative framework, systematically evaluating its strengths and weaknesses against these widely used baselines.
In addition, the design implications and guidelines we derive are constrained by the scope of this study. They are based on a single AR puppeteer system, a single pick-and-place teleoperation task, and two specific modality configurations (GO and VG) with a particular participant sample. Consequently, they should be used as actionable recommendations for similar AR-based robotic teleoperation scenarios, rather than definitive or universal rules for multimodal HRI.
@@
From a design perspective, future work should address the limitations of both input modalities. The voice command techniques could benefit from a more robust and adaptive recognition engine, potentially fine-tuned to domain-specific vocabularies and augmented with real-time clarification mechanisms for ambiguous queries. For gesture input, improvements in gesture-to-action mappings, more resilient tracking algorithms, and clearer visual or haptic feedback could mitigate usability issues and reduce interruptions during rapid movement. Beyond improving the underlying technology, future studies should explore richer task designs that engage voice and gesture across multiple stages and roles, examining different modality combinations and division of labor over time. Finally, expanding evaluation beyond a controlled lab setting and diversifying task scenarios will be essential for assessing the generalizability of AR-based puppeteering to real-world HRI applications.
7 Conclusion
@@
This study explored the impact of gesture-only versus multimodal (VG) interactions in an AR-based robot puppeteering system. Results showed that GO consistently outperformed VG in task performance, usability, and user satisfaction. Voice input benefited novice users, narrowing the performance gap with experts; however, it introduced cognitive overhead and system inefficiencies that affected overall UX. As an early-stage exploratory study, our findings highlight the importance of optimizing modality integration and suggest that hand gesture-based interaction remains a robust solution for robot puppeteer metaphoric system in AR – especially for domain experts – while multimodal designs should be tailored to users’ expertise. Future work should explore adaptive, user-aware multimodal systems that balance flexibility with reliability across diverse user contexts.
[…]
Acknowledgements
Acknowledgements are not compulsory. Where included they should be brief. Grant or contribution numbers may be acknowledged.
Please refer to Journal-level guidance for any specific requirements.
Declarations
  • Funding
  • Conflict of interest/Competing interests (check journal-specific guidelines for which heading to use)
  • Ethics approval and consent to participate
  • Consent for publication
  • Data availability
  • Materials availability
  • Code availability
  • Author contribution